A coin slides over a frictionless plane and across an xy coordinate system from the origin to a point with xy coordinates (3.77 m, 4.84 m) while a constant force acts on it. The force has magnitude 1.50 N and is directed at a counterclockwise angle of 100° from the positive direction of the x axis. How much work is done by the force on the coin during the displacement?
Answer #1

Distance, d = sqrt(3.77^2 + 4.84^2) = 6.14 m

Angle, phi = arctan (4.84/3.77) = 52.08°

Work, W = F d cos(theta)

W = 1.50 x 6.14 x cos(100 - 52.08)

W = 6.17 J

Comment in case any doubt please rate my answer...